OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. 2 B. 3/4 C. -3/4 D. -2

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{-7-3}{4--1}=slope\] \[\LARGE \frac{-7-3}{4+1}=slope\]

OpenStudy (anonymous):

-7-3/4+1=-10/5

pooja195 (pooja195):

good what is -10 divided by 5 ?

OpenStudy (anonymous):

so my answer is -2

pooja195 (pooja195):

Correct! Well done =)

OpenStudy (anonymous):

Thank you!

pooja195 (pooja195):

yw ^-^

OpenStudy (anonymous):

could you help me on 2 more

pooja195 (pooja195):

Sure =)

OpenStudy (anonymous):

Im gonna close this question

OpenStudy (anonymous):

I'll tag you in my new one

pooja195 (pooja195):

Ok :)